# LiDAR Depth Map Guided Image Compression Model

### v2

@@ 2 Examined Scenario
As a first case, we consider to have the LiDAR map available in its original uncompressed format. Thus, we examine the situation where the RGB image and the corresponding LiDAR map are simultaneously acquired from the device, and the uncompressed LiDAR map is stored without compression on the acquisition device, as it will be utilized for the various imaging manipulations previously mentioned. Then, the uncompressed LiDAR data will assist in both encoding the RGB image, which will be then locally stored on the device, and decoding the reconstructed version. In essence, this approach capitalizes on the presence of the LiDAR map on the device to enhance the compression of the RGB image. In this context, we will not consider the rate cost of the LiDAR map, as its presence on the device is independent of the RGB image compression goal.
@@ 4.2 Results
Furthermore, the red curve corresponds to the performance of our model when considering the case in which the uncompressed LiDAR data is unavailable, such as during transmission. In this context, the LiDAR information needs to be compressed separately. In our experiments, it undergoes separate compression using our model, where we specifically provide zero matrices as input to the prompts $l_{a}$ and $l_{s}$ and set $m_{\lambda}=1$ to maximize the quality of the compressed LiDAR map. Then, the compressed LiDAR map is used for assisting the RGB image encoding and decoding. Of course, the final bitstream will consist of both the latent spaces of the RGB image and the LiDAR map. Interestingly, the resulting performance is comparable to that achieved by the model using the original LiDAR map, both in terms of PSNR and SSIM. In fact, the increase in coding cost while encoding the LiDAR depth map is only marginal (around 9% on average), even due to the significantly lower spatial resolution of the LiDAR map in comparison to the RGB image. Moreover, using the compressed LiDAR map in place of the original one does not significantly affect the compression process. Specifically, in terms of Bjøntegaard’s metric, integrating compressed LiDAR results in an average PSNR gain of $0.57$ dB and an average bit rate reduction of $10\%$ compared to its absence (see Tab. 1 for recap). Notably, we remark that we have not retrained the model using the compressed LiDAR maps.
@@ 5 Conclusion
In this paper, we have shown that integrating the LiDAR depth map within the compression process of an RGB camera image results in substantial performance improvements. As the use of LiDAR in mobile devices continues to grow, our work can serve as a starting point for a new research direction. There are two aspects that deserve further exploration in our study. First, the expansion of the experiments to include additional and more variegated RGB-LiDAR datasets, which are expected to increase considering the growing prevalence of LiDAR technology. Secondly, while our findings demonstrate that utilizing a decoded LiDAR map at the decoder side does not compromise the performance compared to the model using the original map, we believe that employing a combined compression approach for both the image and LiDAR map could potentially yield even greater enhancements. We leave this possibility for exploration in future research endeavors.

### v3

@@ 2 Examined Scenario
As a first case, we consider to have the LiDAR map available in its original uncompressed format. Thus, we examine the situation where the RGB image and the corresponding LiDAR map are simultaneously acquired from the device, and the uncompressed LiDAR map is stored without compression on the acquisition device, as it will be utilized for the various imaging manipulations previously mentioned. Then, the uncompressed LiDAR data will assist in both encoding the RGB image, which will be then locally stored on the device, and decoding the reconstructed version. In essence, this approach capitalizes on the presence of the LiDAR map on the device to enhance the compression of the RGB image. In this context, we will not consider the rate cost of the LiDAR map, as its presence on the device is independent of the RGB image compression goal. Of course, this initial case is confined to an encoding and decoding process performed locally on the device.
@@ 4.2 Results
Furthermore, the red curve corresponds to the performance of our model when considering the case in which the uncompressed LiDAR data is unavailable, such as during transmission. In this context, the LiDAR information needs to be compressed separately. In our experiments, it undergoes separate compression using our model, where we specifically provide zero matrices as input to the prompts $l_{a}$ and $l_{s}$ and set $m_{\lambda}=1$ to maximize the quality of the compressed LiDAR map. Then, the compressed LiDAR map is used for assisting the RGB image encoding and decoding. Of course, the final bitstream will consist of both the latent spaces of the RGB image and the LiDAR map. Interestingly, the resulting performance is comparable to that achieved by the model using the original LiDAR map, both in terms of PSNR and SSIM. In fact, the increase in coding cost when encoding the LiDAR depth map is only marginal. On average, it accounts for about 9% of the total cost, which includes both the RGB image and the LiDAR map. This is due to the significantly lower spatial resolution of the LiDAR map compared to the RGB image. Moreover, using the compressed LiDAR map in place of the original one does not significantly affect the compression process. Specifically, in terms of Bjøntegaard’s metric, integrating compressed LiDAR results in an average PSNR gain of $0.57$ dB and an average bit rate reduction of $10\%$ compared to its absence (see Tab. 1 for recap). Notably, we remark that we have not retrained the model using the compressed LiDAR maps.
@@ 5 Conclusion
In this paper, we have shown that integrating the LiDAR depth map within the compression process of an RGB camera image results in substantial performance improvements. As the use of LiDAR in mobile devices continues to grow, our work can serve as a starting point for a new research direction. There are some aspects that deserve further exploration in our study: (i) extending the experiments to include additional and more variegated RGB-LiDAR datasets, which are expected to increase considering the growing prevalence of LiDAR technology; (ii) investigating the universality of the method by integrating the LiDAR map into various image compression frameworks and assessing its efficiency; (iii) while our findings demonstrate that utilizing a decoded LiDAR map at the decoder side does not compromise the performance compared to the model using the original map, we believe that employing a combined compression approach for both the image and LiDAR map could potentially yield even greater enhancements. We leave this possibility for exploration in future research endeavors.